What mass of potassium nitrate, KNO3, is contained in 400 cm3

of an aqueous

solution of KNO3 that has a concentration of 50.0 ppm K

+

. (Assume the only source of K+

in the

solution is KNO3). Relative Atomic Mass Data: K -39.1, N-14.0, O-16.0



1
Expert's answer
2021-11-05T00:55:55-0400

Solution:

Convert ppm to mg/L:

1 ppm = 1 mg L-1

Hence,

(50.0 ppm) × (1 mg L-1 / 1 ppm) = 50.0 mg L-1 K+


The molar mass of K+ is 39.1 g mol-1

Volume of an aqueous solution is 400 cm3 or 0.40 L

Hence,

(50.0 mg K+ / 1 L) × (1 g / 1000 mg) × (1 mol K+ / 39.1 g K+) × (0.40 L) = 0.0005115 mol K+

KNO3 → K+ + NO3-

According to stoichiometry:

Moles of KNO3 = Moles of K+ = 0.0005115 mol


Calculate the mass of KNO3:

The molar mass of KNO3 is 101.1g mol-1

Hence,

(0.0005115 mol KNO3) × (101.1 g KNO3 / 1 mol KNO3) = 0.0517 g KNO3 = 51.7 mg KNO3


Answer: The mass of potassium nitrate (KNO3) is 51.7 mg